Ms. morrision learned that $28 was withdrawn from her checking account when she purchased a scarf in Paris for 126 francs. about how much money would have been withdrawn from her account if she spent 172 francs at a restuarant?
Found 2 solutions by rfer, jorel1380:
Answer by rfer(16322)

28/126=x/172
126x=172*28
x=4816/126
x=38.22

Answer by jorel1380(3719)

28/126=x/172
126x=4816
x=38.22..